# ALMA chemical survey of disk-outflow sources in Taurus (ALMA-DOT)

### v2

@@ Appendix A Atlas of velocity channel maps
The base of the red-shifted wind (with reference to the SE side with respect to the axis, as the NW side hosts the emission from the receding part of the disk) appears resolved in the channel maps at low velocities between VLSR= 7.4 and 9.4 km s. Although the emission pattern is not clear, this velocity interval is consistent with the one expected under the DW hypothesis. In this scenario, the flow starts with at the footpoint (corresponding to VLSR= Vsys), and it is gradually faster as it leaves the disk. However, due to the absorption of the ambient medium, no material with VLSR 7.3 km s is visible (see, e.g. Fig. 5). On the other hand, the maximum VLSR at which the wind base will be imaged is VLSR= Vsys+ , where is the poloidal velocity achieved at an altitude 60 au above the disk. Compared with the curves of the DW model discussed in Section 5.4, is expected to be between 2.0 and 3.5 km sfrom B2 to B1a (see Fig. 13, where the observed quantities are normalized by the corresponding and ). Projecting along the l.o.s. and adding Vsys, one finds that the base of the flow should be visible in channel maps up to VLSR 8.4 – 9.5 km s, as we find, with the highest VLSR corresponding to inner B1a. As predicted by the model, in the channel maps at VLSR km sthe wind base is not resolved.
@@ Appendix B Atlas of transverse Position-Velocity diagrams and application of the ellipse fitting
Figures 18, 19 and 20 present an atlas of the transverse Position-Velocity diagrams (PV⟂) obtained for the red lobe with a 03-wide virtual slit centered on the axis and oriented at PA=318∘ (i.e. perpendicular to the axis and with positive offsets toward NW), at separations from the source from 35 to 10″ in steps of 05
@@
The critical aspect of the procedure is the difficulty in identifying unambiguously the traces corresponding to each shells B1a, B1b, and B2 in the faint or crowded regions of the transverse PV diagrams. To give an identification as reliable as possible, a cross-check on the positions of the emission peaks common to different projections of the datacube was performed. Such a correspondence is expected if the observed features originate from cuts of coherent structures in the datacube. To this aim, we first found sequences of relative intensity peaks in each diagram that could trace a feature, all above an emission threshold of a 3 level, where mJy beam-1 is the average over all panels estimated in a region along the axis at VLSR 20 km s, where the signal of the traces is almost absent. We then looked for the signature of these sequences in the other two projections of the data cube. In particular, the position of the traces in the fan of the PV∥ diagram (Fig. 5) identifies the expected velocity of each feature along the axis, while in the channel maps, the position of the peaks at the intersection of the arc-like features with the virtual perpendicular cut at a distance from the source provides information on the expected offset of the emission in the PV⟂ diagram of offset at the velocity of the channel map. The quantitative comparison of the coordinates in the three projections allowed the selection of the sequence of peaks belonging to a single feature in each diagram.
@@
Then, a mathematical fitting routine based on a least-squares algorithm was applied to each identified sequence of relative intensity peaks. The fitting is anchored to points in the upper arc and side edges, but not in the lower arc, as the latter is difficult to identify due to overlap of low-velocity traces, signal absorption, and deformation due to flow inclination. The appearance of the traces in the PV⟂ projections suggests an ellipse as the geometrical form closest to the actual shape of the features, so the fit requires a number of input points greater than five. Sampling sets that met the above criteria could be found for shells B1a and B1b at offsets from 45 to 7″from the star, and for B2 from 6″to 10″. The input points are reported as crosses in Figures 18 - 20, with the fitted ellipses drawn in the corresponding color. The routine provides the coordinates of the center, the length of the semi-major and semi-minor axes, and the inclination with respect to the horizontal direction (i.e. VLSR= constant).
@@
The SE and NW edges of the fitted ellipses, that is, the points on the curve farthest from the axis, are considered the vertices of each feature at the borders of the shell (in a couple of cases where the geometric extreme of the ellipse had no emission associated, the closest emission peak was selected). The coordinates of these points provide the quantities necessary for the tomographic reconstruction.
@@
With respect to the error estimate, statistically, the applied fitting procedure is expected to reduce the uncertainty with respect to that of the input points. In practice, however, the advantage is contrasted by the difficulty in identifying the traces, the lack of reliable sampling points in the lower part of the curve, the difference of the actual trace shape with respect to a geometrical ellipse, and the absence of emission at the extremes in a few cases. Therefore, we consider the uncertainty in the final coordinate values comparable to that of the input points in the spectral images, that is, in position and km s in velocity, reported as a small magenta ellipse in the right panels.
@@ Appendix D Estimate of the radial velocity V
In Section 4.1 the assumption is made that the wind flows along the detected shells. Formally, this translates into the fact that in the cylindrical coordinate system aligned with the red-shifted flow axis and centered on the star, the poloidal velocity is tangent to the shell surface. To test the validity of this assumption, one should verify that ), where is the local tangent angle to the shell, determined from the derivative of the measured and reported in Footnote 3. While and at a given offset from the source are found directly from the measurements at the lateral apexes of the fitted ellipses in one single diagram, the search for is complicated by the combination of the detected variation of along the shell and the inclination of the flow axis. With reference to Fig. 21, and assuming axisymmetry, to determine the radial velocity corresponding to the axial velocity evaluated at offset , one has to measure VLSR at either point A or point B of the same shell section (the blue circle in the figure), that is, at the points where the projection of along the l.o.s. vanishes:
However, points A and B fall at offsets and . Unfortunately, with the given geometry, these points correspond in almost all cases to regions of the diagrams with faint or confused emission, out of the field of view, hidden by the brightness of the gas close to the source or lost because of the CO auto-absorption around Vsys. These factors make the determination of extremely difficult.
@@
Appendix E Comparison with the tomographic procedure in de Valon et al. (2022).
A tomographic analysis similar to the one presented in Section 4.1 was applied in de Valon et al. (2022) to derive the morphology and the kinematic properties of the molecular outflow from the young star DG Tau B. In this source the flow presents a complex system of internal substructures with characteristics analogous to those of the flow from HL Tau. However, the adopted technique is not identical and leads to different outcomes.
The system of Eq. 1 is equivalent to the one in Eqs. (4)-(7) in Sect. 3 of de Valon et al. (2022)777the inclination angle of the flow with respect to the line of sight differs by 180°with respect to ours, which introduces a minus sign in . In both cases, and are derived from the values of VLSR extracted along the bell profile in the PV⟂ diagrams. The difference resides in the selection of the relevant points, a choice that derives from a different assumption regarding the distribution of the gas in the flow. In de Valon et al. (2022) the wind is considered homogeneous in the transverse direction to the axis. Thus, the points are selected with continuity throughout the bell profile. Under the assumption of a perfectly symmetric flow with respect to the axis, the procedure gives at each distance from the source two continuous functions =VLSR and =VLSR for varying offset . From these, maps of and in the plane are derived over a transversely homogeneous wind (see Fig. 3 in de Valon et al. (2022)). The presence of the observed arcs and cusps in the channel maps and of the ellipse-shaped traces in the PV diagrams does not enter the analysis at this stage. In fact, the features are interpreted as emission enhancements moving in a homogeneous flow. The justification of such enhancements invokes the presence of perturbations in the flow caused by a non-stationary mechanism, influencing either the accretion flow through the disk or the ejection of the axial jet (see Fig. 19 in de Valon et al. (2022)).
@@ Appendix F Evaluation of the dynamical time of the substructures
In the top panel of Fig. 22 we report the dynamical time estimated as , as a function of the distance from the source , for the three substructures analyzed. The dynamical time is higher for the slower outer shells, and for each shell increases monotonically with distance from the source, from 4.8 to 6 102 years for B1a and from 6.0 to 7.1 102 years for B1b, over a distance from 0.9 to 1.4 103 au from the source. For the outer shell B2 is measured to vary between 7.6 and 10 102 years over a distance of 1.2 to 1.9 103 au from the source. The observed trends are compatible with a steady ejection of material and steady propagation along the nested shells as postulated in a magnetohydrodynamic disk wind. The applicability of the disk wind scenario is quantitatively confirmed by Fig. 22, bottom panel, in which the dynamical times derived for the three shells, normalized through their footpoint radii and corresponding Keplerian velocities, are compared to the analogous quantity estimated for the DW models L2 and N5 (cf. Section 5.4). The points show the same trend as the curves, closely following the one for model N5.
@@
Appendix G Comparison with models - variation of along the streamline
Figure 23 illustrates the similarity of the trend of variation of the specific angular momentum along a streamline. Models L2 (Lesur 2021) and N5 (Zimniak et al. 2024) are calculated for =1 au and a star of one solar mass. The points for shell B2 are normalized by the radius of the retrieved footpoint and by the corresponding Keplerian velocity for a stellar mass of 2.1 M⊙. In this way the wind streamline originates in all cases at 1 au, where it is usual to assume that the turbulence starts, so as to have the most effective comparison between the turbulence- and AD-dominated disk wind models (N and L types of solutions, respectively). In all cases the values are seen to follow closely the curve at constant =1, with a progression toward higher with distance, as expected in steady MHD disk winds.

### v3

@@ Acknowledgements.
The base of the redshifted wind (with reference to the SE side with respect to the axis, as the NW side hosts the emission from the receding part of the disk) appears resolved in the channel maps at low velocities between VLSR= 7.4 and 9.4 km s. Although the emission pattern is not clear, this velocity interval is consistent with the one expected under the DW hypothesis. In this scenario, the flow starts with at the footpoint (corresponding to VLSR= Vsys), and it is gradually faster as it leaves the disk. However, due to the absorption of the ambient medium, no material with VLSR 7.3 km s is visible (see, e.g., Fig. 5). On the other hand, the maximum VLSR at which the wind base will be imaged is VLSR= Vsys+ , where is the poloidal velocity achieved at an altitude 60 au above the disk, where au is the spatial resolution element. Compared with the curves of the DW model discussed in Section 5.4, is expected to be between 2.0 and 3.5 km sfrom B2 to B1a (see Fig. 13, where the observed quantities are normalized by the corresponding and ). Projecting along the l.o.s. and adding Vsys, one finds that the base of the flow should be visible in channel maps up to VLSR 8.4 – 9.5 km s, as we find, with the highest VLSR corresponding to inner B1a. As predicted by the model, in the channel maps at VLSR km sthe wind base is not resolved.
@@ Appendix B Atlas of transverse position-velocity diagrams and application of the ellipse fitting
Figures 19, 20 and 21 present an atlas of the transverse Position-Velocity diagrams (PV⟂) obtained for the red lobe with a 03-wide virtual slit centered on the axis and oriented at PA=318∘ (i.e., perpendicular to the axis and with positive offsets toward NW), at separations from the source from 35 to 10″ in steps of 05, as indicated by the horizontal dashed lines in Fig. 18.
@@
The critical aspect of the procedure is the difficulty in identifying unambiguously the traces corresponding to each shells B1a, B1b, and B2 in the faint or crowded regions of the transverse PV diagrams. To give an identification as reliable as possible, a cross-check on the positions of the emission peaks common to different projections of the datacube was performed. Such a correspondence is expected if the observed features originate from cuts of coherent structures in the datacube. To this aim, we first found sequences of relative intensity peaks in each PV⟂ diagram that could trace a feature, all above an emission threshold of a 3 level, where mJy beam-1 is the average over all panels estimated in a region along the axis at VLSR 20 km s, where the signal of the traces is almost absent. We then looked for the signature of these sequences in the other two projections of the datacube. In particular, the position of the traces in the fan of the PV∥ diagram (Fig. 5) identifies the expected velocity of each feature along the axis. On the other hand, the position of the traces in the channel maps at each offset gives information on the expected offset of the emission in the PV⟂ diagram formed at the distance from the source, and at the velocity of the selected channel map. We verified this correspondence by reporting on the channel map at a given velocity the positions of the traces in the PV⟂ diagrams at each offset , at the selected velocity. An example is given in Fig. 18 for the map at VLSR= 16.2 km s. The quantitative comparison of the coordinates in the three projections allowed the selection of the sequence of peaks belonging to a single feature in each diagram.
@@
Then, a mathematical fitting routine based on a least-squares algorithm was applied to each identified sequence of relative intensity peaks. The fitting is anchored to points in the upper arc and side edges, but not in the lower arc, as the latter is difficult to identify due to overlap of low-velocity traces, signal absorption, and deformation due to flow inclination. The appearance of the traces in the PV⟂ projections suggests an ellipse as the geometrical form closest to the actual shape of the features, so the fit requires a number of input points greater than five. Sampling sets that met the above criteria could be found for shells B1a and B1b at offsets from 45 to 7″from the star, and for B2 from 6″to 10″. The input points are reported as crosses in Figures 19 - 21, with the fitted ellipses drawn in the corresponding color. The routine provides the coordinates of the center, the length of the semi-major and semi-minor axes, and the inclination with respect to the horizontal direction (i.e., VLSR= constant).
@@
The SE and NW edges of the fitted ellipses, that is, the points on the curve farthest from the axis, were considered the vertices of each feature at the borders of the shell (in a couple of cases where the geometric extreme of the ellipse had no emission associated, the closest emission peak was selected). The coordinates of these points provided the quantities necessary for the tomographic reconstruction.
@@
With respect to the error estimate, statistically, the applied fitting procedure is expected to reduce the uncertainty with respect to that of the input points. In practice, however, the advantage is contrasted by the difficulty in identifying the traces, the lack of reliable sampling points in the lower part of the curve, the difference of the actual trace shape with respect to a geometrical ellipse, and the absence of emission at the extremes in a few cases. Therefore, we considered the uncertainty in the final coordinate values to be comparable to that of the input points in the spectral images, that is, in position and km s in velocity, reported as a small magenta ellipse in the right panels.
@@ Appendix D Estimate of the radial velocity V
In Section 4.1 the assumption is made that the wind flows along the detected shells. Formally, this translates into the fact that in the cylindrical coordinate system aligned with the redshifted flow axis and centered on the star, the poloidal velocity is tangent to the shell surface. To test the validity of this assumption, one should verify that ), where is the local tangent angle to the shell, determined from the derivative of the measured and reported in Footnote 3. While and at a given offset from the source are found directly from the measurements at the lateral apexes of the fitted ellipses in one single diagram, the search for is complicated by the combination of the detected variation of along the shell and the inclination of the flow axis. With reference to Fig. 22, and assuming axisymmetry, to determine the radial velocity corresponding to the axial velocity evaluated at offset , one has to measure VLSR at either point A or point B of the same shell section (the blue circle in the figure), that is, at the points where the projection of along the l.o.s. vanishes:
However, points A and B fall at offsets and . Unfortunately, with the given geometry, these points correspond in almost all cases to regions of the diagrams with faint or confused emission, out of the field of view, hidden by the brightness of the gas close to the source, or lost because of the CO auto-absorption around Vsys. These factors make the determination of extremely difficult.
@@
Appendix E Comparison with the tomographic procedure in DeValon22
A tomographic analysis similar to the one presented in Section 4.1 was applied in DeValon22 to derive the morphology and the kinematic properties of the molecular outflow from the young star DG Tau B. In this source the flow presents a complex system of internal substructures with characteristics analogous to those of the flow from HL Tau. However, the adopted technique is not identical and leads to different outcomes.
The system of Eq. 1 is equivalent to the one in Eqs. (4)-(7) in Sect. 3 of DeValon22.777the inclination angle of the flow with respect to the line of sight differs by 180°with respect to ours, which introduces a minus sign in In both cases, and are derived from the values of VLSR extracted along the bell profile in the PV⟂ diagrams. The difference resides in the selection of the relevant points, a choice that derives from a different assumption regarding the distribution of the gas in the flow. In DeValon22 the wind is considered homogeneous in the transverse direction to the axis. Thus, the points are selected with continuity throughout the bell profile. Under the assumption of a perfectly symmetric flow with respect to the axis, the procedure gives at each distance from the source two continuous functions =VLSR and =VLSR for varying offset . From these, maps of and in the plane are derived over a transversely homogeneous wind (see Fig. 3 in DeValon22). The presence of the observed arcs and cusps in the channel maps and of the ellipse-shaped traces in the PV diagrams does not enter the analysis at this stage. In fact, the features are interpreted as increases in density moving in a homogeneous flow. The justification of such increases invokes the presence of perturbations in the flow caused by a non-stationary mechanism, influencing either the accretion flow through the disk or the ejection of the axial jet (see Fig. 19 in DeValon22).
@@ Appendix F Evaluation of the dynamical time of the substructures
In the upper panel of Fig. 23 we report the dynamical time estimated as , as a function of the distance from the source , for the three substructures analyzed. The dynamical time is higher for the slower outer shells, and for each shell increases monotonically with distance from the source, from 4.8 to 6 102 years for B1a and from 6.0 to 7.1 102 years for B1b, over a distance from 0.9 to 1.4 103 au from the source. For the outer shell B2 is measured to vary between 7.6 and 10 102 years over a distance of 1.2 to 1.9 103 au from the source. The observed trends are compatible with a steady ejection of material and steady propagation along the nested shells as postulated in a magnetohydrodynamic disk wind. The applicability of the DW scenario is quantitatively confirmed by Fig. 23, bottom panel, in which the dynamical times derived for the three shells, normalized through their footpoint radii and corresponding Keplerian velocities, are compared with the analogous quantity estimated for the DW models L2 and N5 (cf. Section 5.4). The points show the same trend as the curves, closely following the one for model N5.
@@
Appendix G Comparison with models: Variation of along the streamline
Figure 24 illustrates the similarity of the trend of variation of the specific angular momentum along a streamline in the DW models examined and in the observed wind. Models L2 (Lesur21) and N5 (Zimniak24) are calculated for =1 au and a star of one solar mass. The points derived for shell B2 from the data are normalized by the radius of the retrieved footpoint and by the corresponding Keplerian velocity, for a stellar mass of 2.1 M⊙. In this way the wind streamline is considered to originate at 1 au, which is usually assumed to be the upper limit of the extension of the turbulent inner disk. In this way we have the most effective comparison between the turbulence- and AD-dominated disk wind models (N and L types of solutions, respectively) and the observed wind. In all cases the values are seen to follow closely the curve at constant =1, with a progression toward higher with distance, as expected in steady MHD disk winds.